OpenStudy (anonymous):

3x/2+5/24=5x/8+13/12

OpenStudy (ahsome):

Rewrite equation: \[\frac{3x}{2}+\frac{5}{24}=\frac{5x}{8}+\frac{13}{12}\] Change all so thet have a common denominator (In this case \(24\)) \[\frac{3x*12}{2*12}+\frac{5}{24}=\frac{5x*3}{8*3}+\frac{13*2}{12*2}\] Solve the multiplication: \[\frac{36x}{24}+\frac{5}{24}=\frac{15x}{24}+\frac{26}{24}\] Times both sides by 24 to remove the denominator \[24(\frac{36x}{24}+\frac{5}{24})=24(\frac{15x}{24}+\frac{26}{24})\] Simplify: \[36x+5=15x+26\] Put common terms together \[36x-15x=26-5\] Simplify, then solve: \[21x=21\] \[x=\frac{21}{21}\] \[x=1\] Hope that helps, @Laura161. Just say if you don't understand. If this helped you, please click the \(Best-Response\) and \(Close-Question\) button :)